# RobGC: Towards Robust Graph Condensation

### v1

@@ IV-A Experimental Settings
Dataset ($r$) Cora (5.80%) Citeseer (3.30%)
Noise level (%) 0 20 40 60 80 100 0 20 40 60 80 100
Whole 86.77±0.29 82.90±0.37 78.77±0.33 76.43±0.12 74.00±0.49 70.90±1.13 77.70±0.37 75.20±0.22 72.80±0.14 70.60±0.29 69.53±0.29 69.17±0.26
GCond Plain 80.50±0.08 77.37±0.24 73.97±0.26 69.10±0.78 68.70±0.50 64.67±0.79 75.47±0.26 70.70±1.16 70.13±0.19 69.40±0.24 67.13±0.54 66.23±0.12
Jaccard 79.87±1.44 75.90±0.08 74.23±0.59 71.87±0.31 71.73±0.24 67.47±0.52 74.03±0.21 71.80±0.49 71.57±0.33 70.60±0.75 68.07±0.25 68.00±0.29
SVD 77.80±0.14 76.90±0.28 74.30±0.28 71.03±0.05 70.87±0.41 68.47±0.45 74.30±0.22 71.97±0.19 70.37±0.34 70.10±0.08 68.63±0.42 68.27±0.21
kNN 78.43±0.17 77.80±0.37 74.63±0.45 70.87±0.31 69.57±0.09 65.73±0.73 74.30±0.08 71.63±0.69 71.20±0.96 70.13±0.17 68.27±0.24 67.30±0.22
Ours 80.90±0.51 78.93±0.66 76.30±0.29 75.53±0.17 73.00±0.22 70.57±0.21 75.87±0.41 75.47±0.05 75.13±0.21 74.80±0.22 73.83±0.29 72.90±0.08
GCDM Plain 77.73±0.70 67.97±1.15 63.47±0.90 61.87±0.17 56.70±0.43 53.87±1.30 74.70±0.14 70.90±0.37 69.77±0.21 66.10±0.57 65.03±1.02 60.47±0.39
Jaccard 77.57±0.46 72.10±0.16 68.47±0.24 65.30±0.43 61.77±0.69 61.10±0.79 74.53±0.05 72.73±0.17 70.57±0.39 67.40±0.29 66.63±0.98 64.73±0.42
SVD 76.43±0.31 73.30±0.22 71.50±0.22 63.37±0.46 62.40±0.82 57.63±2.67 74.60±0.29 71.23±0.21 70.37±0.46 67.37±0.19 66.30±0.43 62.67±0.40
kNN 76.23±0.54 68.60±0.37 64.87±1.26 62.73±0.37 58.43±0.33 55.30±1.20 74.07±0.17 72.23±0.47 70.03±0.21 69.73±0.21 66.07±0.09 61.30±0.16
Ours 77.85±0.29 74.13±0.21 72.43±0.25 67.77±0.53 63.87±0.26 63.77±0.60 74.87±0.12 73.67±0.05 71.87±0.09 71.60±0.16 70.77±0.05 70.17±0.08
GCEM Plain 77.77±0.33 71.80±0.22 68.10±0.43 62.67±0.25 59.13±0.50 56.97±0.46 74.17±0.56 72.00±0.08 68.50±0.37 67.27±0.50 65.20±0.29 61.47±0.59
Jaccard 77.80±0.16 73.67±0.21 71.17±0.56 65.47±0.19 63.30±0.33 62.63±0.33 74.45±0.16 72.57±0.19 71.80±0.08 64.47±0.17 65.40±0.22 67.33±0.57
SVD 75.03±0.26 72.07±0.91 69.17±0.25 64.47±0.74 63.73±0.73 63.37±0.26 73.40±0.36 72.60±0.08 68.87±0.05 68.53±0.17 67.33±0.37 66.43±0.25
kNN 77.07±0.48 72.57±0.31 69.97±0.17 66.93±0.33 65.93±0.21 63.00±0.36 73.20±0.37 70.10±0.29 68.57±0.31 67.97±0.25 66.97±0.33 65.20±0.22
Ours 77.84±0.21 74.10±0.43 72.83±0.46 68.97±0.39 66.10±0.22 63.70±0.33 74.47±0.24 72.67±0.33 72.10±0.08 71.30±0.24 71.07±0.17 68.07±0.17
Dataset ($r$) Pubmed (0.17%) Ogbn-arxiv (0.50%)
Noise level (%) 0 20 40 60 80 100 0 20 40 60 80 100
Whole 88.00±0.08 85.97±0.25 81.90±0.28 80.60±0.08 78.37±0.25 77.50±0.22 69.94±0.05 66.42±0.10 64.64±0.27 63.47±0.19 62.21±0.07 61.24±0.08
GCond Plain 81.40±0.14 78.10±0.22 76.20±0.08 75.10±0.57 71.83±0.12 69.77±0.29 59.29±0.14 53.67±0.85 52.22±0.45 52.01±0.72 48.38±0.15 46.43±0.45
Jaccard 81.00±0.37 78.27±0.21 75.83±0.25 74.97±0.09 73.83±0.09 73.67±0.56 58.20±0.24 54.12±0.41 53.06±0.19 52.13±0.26 49.56±0.47 47.80±0.21
SVD 80.27±0.60 78.60±0.16 78.17±0.05 77.60±0.16 76.63±0.05 75.53±0.69 58.29±0.18 53.49±0.40 53.12±0.49 52.33±0.95 48.86±0.54 47.97±0.75
kNN 80.63±0.21 80.40±0.24 78.80±0.28 76.97±0.05 75.37±0.26 73.63±0.85 58.96±0.69 54.18±0.73 53.02±0.85 52.24±1.56 48.61±2.27 47.13±1.42
Ours 81.53±0.09 80.77±0.12 80.73±0.74 79.47±0.40 77.00±0.24 76.80±0.28 59.60±0.22 54.97±0.54 54.01±0.34 53.83±0.52 50.00±0.42 49.67±0.26
GCDM Plain 83.37±0.38 76.67±0.57 74.90±0.29 72.80±1.13 72.70±0.36 70.73±0.19 57.77±0.33 52.20±0.43 52.02±0.16 49.99±0.61 49.10±0.45 46.24±0.45
Jaccard 82.33±0.17 78.40±0.08 75.93±0.12 74.20±0.22 72.23±0.05 72.60±0.08 58.68±0.18 54.99±0.27 53.28±0.52 50.96±0.23 49.86±0.14 48.28±0.82
SVD 79.57±0.61 78.03±0.26 77.57±0.75 77.43±0.17 76.00±0.14 76.40±0.08 57.55±0.49 53.23±0.39 52.42±0.17 50.43±0.16 49.30±0.32 49.22±0.23
kNN 82.17±0.41 79.07±0.45 76.57±0.61 74.77±0.25 72.90±0.14 71.17±0.25 57.44±3.49 52.33±0.17 52.77±0.86 50.75±0.87 49.29±3.05 48.31±0.36
Ours 83.63±0.19 79.23±0.12 78.50±0.16 78.00±0.22 77.70±0.29 77.30±0.43 59.18±0.15 55.38±0.21 53.92±0.31 51.19±0.45 50.38±0.77 49.27±0.36
GCEM Plain 81.97±0.41 72.97±0.45 70.40±0.29 68.67±0.37 65.60±0.14 64.80±0.14 58.87±0.26 55.45±0.47 52.73±0.45 51.66±0.45 51.02±0.26 49.41±0.28
Jaccard 82.01±0.26 77.27±0.05 71.20±0.22 69.37±0.12 67.43±0.05 66.00±0.22 57.34±0.79 56.46±0.15 53.02±0.29 52.12±0.34 51.65±0.21 49.75±0.17
SVD 78.87±0.26 77.83±0.41 75.63±0.05 74.10±0.29 72.90±0.22 71.00±0.24 58.50±0.08 56.97±0.26 53.27±0.42 52.40±0.43 51.70±0.50 49.67±0.79
kNN 79.57±0.71 73.87±0.25 71.00±0.37 69.07±0.68 68.13±1.81 67.93±1.60 57.45±0.48 56.84±0.71 53.75±0.62 52.52±0.98 51.33±1.77 49.08±1.89
Ours 82.03±0.05 79.20±0.08 75.97±0.12 74.43±0.54 73.73±0.21 72.93±0.05 59.52±0.13 57.00±0.21 54.37±0.27 53.87±0.19 52.41±0.64 52.03±0.28
@@
Baselines and Evaluation Settings. We deploy our proposed method on three GC methods with different optimization strategies: (1) GCond [13]: the first GC method that matches the model gradients derived from the training and condensed graph, ensuring the condensed graph effectively preserves the knowledge for GNN training; (2) GCDM [31]: an efficient GC method that minimizes the discrepancy between node distributions of the training and condensed graphs in the feature space. (3) GCEM [32]: a generalized GC method that decomposes the adjacency matrix of the training graph, ensuring that the condensed graph’s eigenvectors align with those selected from the training graph.
@@
Our primary objective is to enhance graph quality rather than merely optimizing model performance. Consequently, we select three state-of-the-art (SOTA) denoising techniques capable of addressing noise in inductive test settings. These methods are deployed with the aforementioned GC methods for comparison: (1) Jaccard [27]: it computes the Jaccard similarity for every pair of connected nodes and discards edges whose similarity is smaller than a specific threshold. (2) SVD [27]: it decomposes the noisy graph through Singular Value Decomposition (SVD) and adopts the low-rank approximation for the denoised graph to reduce the influence of potential high-rank noises. (3) kNN [39]: it identifies the k-nearest neighbors for each node based on their features and incorporates potentially effective edges into the original graph.
@@
Following [27], we assess our method’s resilience to both random noise and adversarial attacks. For random noise, it involves randomly adding fake edges to and removing real edges from the graph. Given the computational intensity of most adversarial attack methods, we utilize the SOTA scalable attack method, PR-BCD [26], to execute global adversarial attacks. We evaluate each method across different noise levels, defined by the ratio of changed edges in the original graph [27]. For random noise, levels range from 0 to 1.0, increasing by 0.2, while for adversarial attacks, levels are set between 0.05 and 0.25, with increments of 0.05. We operate under the premise of uniform noise levels in real-world systems, indicating both training and test graphs share the identical noise level.

### v2

@@ IV-A Experimental Settings
Dataset ($r$) Cora (5.80%) Citeseer (3.30%) Pubmed (0.17%) Ogbn-arxiv (0.50%)
Noise level (%) 0 20 40 100 0 20 40 100 0 20 40 100 0 20 40 100
Whole Dataset 86.8±0.3 82.9±0.4 78.8±0.3 70.9±1.1 77.7±0.4 75.2±0.2 72.8±0.1 69.2±0.3 88.0±0.1 86.0±0.2 81.9±0.3 77.5±0.2 69.9±0.1 66.4±0.1 64.6±0.3 61.2±0.1
GCond Plain 80.5±0.1 77.4±0.2 74.0±0.3 64.7±0.8 75.5±0.3 70.7±1.2 70.1±0.2 66.2±0.1 81.4±0.1 78.1±0.2 76.2±0.1 69.8±0.3 59.3±0.1 53.7±0.8 52.2±0.5 46.4±0.5
Jaccard 79.9±1.4 75.9±0.1 74.2±0.6 67.5±0.5 74.0±0.2 71.8±0.5 71.6±0.3 68.0±0.3 81.0±0.4 78.3±0.2 75.8±0.2 73.7±0.6 58.2±0.2 54.1±0.4 53.1±0.2 47.8±0.2
SVD 77.8±0.1 76.9±0.3 74.3±0.3 68.5±0.5 74.3±0.2 72.0±0.2 70.4±0.3 68.3±0.2 80.3±0.6 78.6±0.2 78.2±0.1 75.5±0.7 58.3±0.2 53.5±0.4 53.1±0.5 48.0±0.8
kNN 78.4±0.2 77.8±0.4 74.6±0.5 65.7±0.7 74.3±0.1 71.6±0.7 71.2±1.0 67.3±0.2 80.6±0.2 80.4±0.2 78.8±0.3 73.6±0.8 59.0±0.7 54.2±0.7 53.0±0.8 47.1±1.4
${\rm STABLE_{G}}$ 80.6±0.2 77.9±0.3 75.0±0.2 69.3±0.2 75.5±0.1 72.8±0.4 72.6±0.1 70.0±0.3 81.3±0.4 79.1±0.2 78.8±0.2 75.7±0.6 59.3±0.2 54.1±0.2 53.3±0.3 48.5±0.1
Ours 80.9±0.5 78.9±0.7 76.3±0.3 70.6±0.2 75.9±0.4 75.5±0.1 75.1±0.2 72.9±0.1 81.5±0.1 80.8±0.1 80.7±0.7 76.8±0.3 59.6±0.2 55.0±0.5 54.0±0.3 49.7±0.3
GCDM Plain 77.7±0.7 68.0±1.1 63.5±0.9 53.9±1.3 74.7±0.1 70.9±0.4 69.8±0.2 60.5±0.4 83.4±0.4 76.7±0.6 74.9±0.3 70.7±0.2 57.8±0.3 52.2±0.4 52.0±0.2 46.2±0.5
Jaccard 77.6±0.5 72.1±0.2 68.5±0.2 61.1±0.8 74.5±0.1 72.7±0.2 70.6±0.4 64.7±0.4 82.3±0.2 78.4±0.1 75.9±0.1 72.6±0.1 58.7±0.2 55.0±0.3 53.3±0.5 48.3±0.8
SVD 76.4±0.3 73.3±0.2 71.5±0.2 57.6±2.7 74.6±0.3 71.2±0.2 70.4±0.5 62.7±0.4 79.6±0.6 78.0±0.3 77.6±0.8 76.4±0.1 57.5±0.5 53.2±0.4 52.4±0.2 49.2±0.2
kNN 76.2±0.5 68.6±0.4 64.9±1.3 55.3±1.2 74.1±0.2 72.2±0.5 70.0±0.2 61.3±0.2 82.2±0.4 79.1±0.5 76.6±0.6 71.2±0.2 57.4±3.5 52.3±0.2 52.8±0.9 48.3±0.4
${\rm STABLE_{G}}$ 77.8±0.4 73.5±0.2 71.7±0.2 62.0±0.7 74.8±0.2 72.9±0.3 71.2±0.2 67.7±0.7 83.4±0.3 78.6±0.4 77.2±0.4 75.6±0.3 58.8±0.3 54.8±0.3 53.4±0.1 49.0±0.3
Ours 77.8±0.3 74.1±0.2 72.4±0.2 63.8±0.6 74.9±0.1 73.7±0.1 71.9±0.1 70.2±0.1 83.6±0.2 79.2±0.1 78.5±0.2 77.3±0.4 59.2±0.1 55.4±0.2 53.9±0.3 49.3±0.4
GDEM Plain 80.9±0.3 78.9±0.3 75.2±0.4 65.1±0.5 75.9±0.6 72.1±0.4 71.5±0.4 67.5±0.5 82.0±0.4 78.8±0.4 77.4±0.3 71.2±0.2 59.9±0.3 55.9±0.4 53.6±0.3 49.1±0.3
Jaccard 81.1±0.3 80.8±0.2 76.3±0.5 68.7±0.4 76.2±0.2 72.6±0.2 71.8±0.1 68.3±0.6 82.2±0.3 80.3±0.1 78.2±0.2 73.0±0.2 58.8±0.8 56.5±0.1 54.0±0.3 49.8±0.2
SVD 79.6±0.3 79.3±0.9 76.1±0.3 69.5±0.3 75.4±0.4 72.8±0.1 71.7±0.1 68.4±0.2 81.2±0.3 80.8±0.4 79.6±0.1 76.0±0.2 59.5±0.1 57.1±0.3 54.3±0.4 49.7±0.8
kNN 81.0±0.5 80.7±0.3 76.2±0.2 69.7±0.4 76.1±0.4 72.4±0.3 71.5±0.3 68.2±0.2 79.7±0.7 79.1±0.2 78.0±0.4 74.9±1.6 59.3±0.5 56.8±0.7 54.8±0.6 49.3±1.9
${\rm STABLE_{G}}$ 81.3±0.2 80.9±0.2 77.4±0.3 68.5±0.3 76.2±0.2 73.2±0.3 72.0±0.3 68.5±0.6 82.1±0.4 80.5±0.1 80.1±0.2 77.4±0.5 60.1±0.9 56.9±0.7 55.1±0.4 50.2±0.3
Ours 82.5±0.2 81.8±0.4 79.9±0.5 70.9±0.2 76.3±0.2 76.0±0.3 75.4±0.1 73.2±0.2 82.2±0.2 81.8±0.2 80.8±0.1 78.9±0.2 60.5±0.1 58.2±0.2 55.8±0.3 52.1±0.3
Dataset ($r$) Flickr (0.5%) Reddit (0.1%) Squirrel (1.7%) Gamers (0.5%)
Noise level (%) 0 20 40 100 0 20 40 100 0 20 40 100 0 20 40 100
Whole Dataset 50.9±0.1 50.2±0.1 49.7±0.1 48.9±0.1 95.3±0.1 93.9±0.0 93.3±0.0 92.1±0.0 44.8±0.5 35.3±0.0 27.5±0.1 24.3±0.2 62.1±0.0 61.9±0.0 61.4±0.1 60.8±0.0
GCond Plain 47.9±0.1 47.5±0.1 47.5±0.1 47.0±0.0 90.7±0.7 89.5±0.9 88.7±0.6 87.0±0.7 36.6±1.5 27.0±1.3 22.4±1.6 21.8±1.5 57.6±0.0 57.4±0.0 56.8±0.5 56.1±0.0
Jaccard 47.7±0.2 47.7±0.5 47.6±0.1 47.1±0.4 90.8±0.5 89.7±0.5 88.9±0.5 87.3±0.5 36.9±1.2 27.3±1.0 24.1±1.4 22.9±1.9 57.7±0.3 57.6±0.1 57.0±0.3 56.4±0.1
SVD 47.8±0.4 47.6±0.1 47.6±0.5 47.2±0.4 90.9±0.8 89.7±0.6 89.4±0.6 87.5±0.6 36.8±1.0 27.5±1.3 24.2±1.5 23.6±1.6 57.9±0.1 57.6±0.2 57.1±0.2 56.4±0.2
kNN 48.1±0.2 47.8±0.2 47.7±0.0 47.4±0.1 90.7±0.5 89.9±0.8 89.3±0.4 87.8±0.4 36.6±1.7 27.8±1.7 24.4±1.7 23.2±1.6 57.8±0.1 57.8±0.1 57.2±0.1 56.7±0.1
${\rm STABLE_{G}}$ 47.9±0.2 47.7±0.1 47.7±0.3 47.2±0.3 90.9±0.4 90.1±0.5 89.5±0.6 87.8±0.5 36.7±1.4 27.7±1.5 24.4±1.4 23.7±1.4 57.9±0.1 57.7±0.1 57.2±0.2 56.6±0.1
Ours 48.4±0.0 48.3±0.0 48.2±0.1 48.0±0.1 91.1±0.5 90.7±0.9 90.5±0.7 89.3±0.7 37.2±1.5 28.3±1.4 25.0±1.2 24.7±1.9 58.4±0.2 58.4±0.1 57.9±0.3 57.8±0.1
GCDM Plain 47.7±0.0 47.2±0.1 47.1±0.2 46.1±0.1 90.8±0.2 89.8±0.4 89.1±0.6 87.1±0.5 36.5±2.7 26.8±1.7 22.3±1.4 21.5±0.9 58.6±0.2 58.1±0.2 57.8±0.1 57.0±0.1
Jaccard 47.9±0.1 47.6±0.0 47.3±0.2 46.4±0.1 91.0±0.3 89.7±0.5 88.9±0.5 87.4±0.1 36.6±2.2 27.2±1.2 23.8±1.2 22.2±1.2 58.7±0.1 58.5±0.1 58.1±0.1 57.2±0.2
SVD 47.7±0.1 47.4±0.0 47.3±0.0 46.3±0.0 90.9±0.4 89.7±0.6 89.4±0.6 87.4±0.5 36.4±2.1 27.6±1.4 23.9±2.3 23.2±1.6 58.7±0.2 58.6±0.1 58.3±0.1 57.6±0.1
kNN 47.9±0.0 47.7±0.3 47.5±0.2 46.4±0.2 91.0±0.2 89.9±0.8 89.3±0.4 88.0±0.3 36.7±1.4 27.6±1.7 23.2±1.4 22.1±0.9 59.0±0.3 58.8±0.1 58.6±0.1 57.7±0.1
${\rm STABLE_{G}}$ 47.8±0.1 47.6±0.0 47.6±0.2 46.4±0.0 91.1±0.3 90.1±0.5 89.5±0.6 87.8±0.3 36.8±1.8 27.8±1.6 24.1±1.1 23.3±1.3 58.8±0.2 58.7±0.2 58.5±0.1 57.7±0.2
Ours 48.3±0.4 48.1±0.3 48.1±0.2 47.3±0.4 91.3±0.5 91.0±0.6 90.6±0.5 89.5±0.7 37.0±1.6 28.2±1.4 24.7±1.8 23.9±0.7 59.4±0.1 59.2±0.1 59.1±0.0 58.6±0.1
GDEM Plain 48.5±0.1 48.0±0.1 47.7±0.1 47.4±0.1 91.7±0.4 91.5±0.3 90.4±0.3 89.4±0.2 37.9±2.2 28.2±1.4 23.6±1.7 22.5±1.2 59.7±0.1 59.4±0.0 59.0±0.5 58.3±0.0
Jaccard 48.4±0.2 48.1±0.1 47.9±0.3 47.5±0.3 91.9±0.1 91.7±0.1 90.7±0.1 89.6±0.4 38.1±1.2 28.6±1.0 25.0±1.1 23.1±1.4 59.7±0.3 59.5±0.1 59.2±0.2 58.4±0.1
SVD 48.5±0.1 48.3±0.0 47.8±0.1 47.6±0.0 91.6±0.2 91.6±0.3 90.7±0.3 89.7±0.1 38.2±1.7 28.8±1.7 25.2±1.3 24.8±1.7 59.8±0.3 59.6±0.1 59.4±0.1 58.5±0.2
kNN 48.6±0.4 48.2±0.1 47.9±0.2 47.6±0.1 91.8±0.5 91.7±0.6 90.8±0.4 89.9±0.2 38.1±1.9 29.0±1.2 25.2±1.2 24.9±1.2 59.9±0.2 59.8±0.1 59.6±0.3 58.6±0.1
${\rm STABLE_{G}}$ 48.6±0.2 48.3±0.2 47.9±0.2 47.7±0.0 91.8±0.3 91.8±0.1 90.8±0.3 90.0±0.3 37.9±1.0 29.4±1.6 25.3±1.1 24.7±1.6 59.9±0.1 59.7±0.2 59.7±0.1 58.6±0.2
Ours 48.9±0.2 48.7±0.2 48.6±0.1 48.5±0.1 92.4±0.1 92.3±0.1 91.7±0.3 91.1±0.2 38.5±1.3 29.8±1.1 25.7±1.1 25.5±1.9 60.3±0.0 60.2±0.1 60.2±0.0 59.5±0.1
TABLE III: The comparison of methods under adversarial attack conditions. Refer to Table II for Noise level=0%.
Dataset ($r$) Cora (5.80%) Citeseer (3.30%) Pubmed (0.17%) Ogbn-arxiv (0.50%)
Noise level (%) 5 10 25 5 10 25 5 10 25 5 10 25
Whole Dataset 84.8±0.1 82.7±0.1 79.1±0.9 76.2±1.1 75.7±0.2 71.6±0.3 83.6±0.1 82.8±0.2 77.6±0.3 65.8±0.2 64.3±0.1 62.2±0.1
GCond Plain 76.8±0.3 74.8±0.7 68.2±0.3 73.3±0.4 71.6±0.3 64.4±0.4 77.6±0.3 75.3±0.2 69.2±0.1 54.4±0.2 51.9±0.2 45.4±0.2
Jaccard 77.1±0.6 76.6±0.1 72.1±0.3 73.9±0.3 71.9±0.5 67.7±0.2 77.4±0.2 76.1±0.1 70.4±0.2 54.8±0.4 52.3±0.2 46.4±0.4
SVD 77.3±0.4 75.8±0.2 70.5±0.6 74.4±0.1 73.7±0.2 69.4±0.4 80.1±0.3 79.9±0.3 78.3±0.3 54.5±2.9 52.6±2.4 47.8±3.0
kNN 77.1±0.9 75.3±0.5 71.6±1.4 74.7±0.1 72.7±0.2 67.0±0.1 78.0±0.2 76.7±0.3 70.8±0.4 54.6±0.8 52.9±0.9 46.0±0.9
${\rm STABLE_{G}}$ 78.0±0.5 76.7±0.4 72.3±0.5 74.6±0.1 73.9±0.3 70.9±0.3 78.8±0.2 77.7±0.2 76.2±0.3 54.7±0.8 52.7±0.3 47.7±0.4
Ours 78.7±1.0 77.8±0.4 73.9±0.2 75.8±0.4 75.4±0.1 73.4±0.5 80.6±0.3 79.9±0.3 78.7±0.4 54.8±0.6 53.1±0.3 48.2±0.4
GCDM Plain 72.7±0.4 71.0±0.3 58.9±0.6 73.1±0.1 71.6±0.1 56.7±0.3 78.1±1.0 75.4±0.4 67.9±0.4 54.5±0.3 48.0±0.9 44.5±0.3
Jaccard 74.7±0.9 73.1±0.6 71.3±0.6 73.6±0.1 72.5±0.2 69.7±0.4 78.8±0.4 76.5±0.3 69.9±0.4 55.0±0.5 50.8±0.2 48.5±0.4
SVD 73.7±0.3 73.1±0.5 65.2±0.2 73.4±0.1 71.9±0.2 67.6±0.1 79.0±0.5 78.6±0.1 76.0±1.2 54.8±0.7 48.7±0.7 46.1±1.6
kNN 73.0±0.3 72.2±1.7 66.9±0.7 73.7±0.7 71.9±0.3 63.2±0.5 78.8±0.4 76.9±0.1 69.8±0.2 55.0±0.2 49.8±0.5 47.2±0.6
${\rm STABLE_{G}}$ 74.9±0.4 74.1±0.7 70.9±0.6 73.7±0.5 72.8±0.4 70.2±0.4 79.1±0.3 78.9±0.2 75.8±0.3 55.2±0.2 51.0±0.4 48.8±0.4
Ours 76.7±0.5 75.4±0.1 71.6±1.2 73.9±0.1 73.4±0.2 71.8±0.1 79.4±0.2 79.2±0.2 77.1±0.1 55.7±0.2 53.8±0.1 50.7±0.3
GDEM Plain 77.2±0.4 76.1±0.4 68.7±0.3 74.4±0.3 72.4±0.2 65.7±0.3 78.1±0.4 76.0±0.3 70.4±0.2 55.6±0.3 52.8±0.6 47.2±0.3
Jaccard 79.3±0.5 77.3±0.3 72.1±0.1 75.3±0.2 73.2±0.3 68.3±0.4 79.1±0.3 76.2±0.2 71.9±0.2 56.8±0.5 53.6±0.5 49.6±0.6
SVD 78.3±0.2 76.7±0.2 70.1±0.2 76.1±0.3 73.3±0.5 69.6±0.3 79.7±0.3 79.5±0.5 78.1±0.6 56.6±0.2 54.0±0.2 49.6±0.3
kNN 78.3±0.3 76.9±0.4 71.2±0.3 75.6±0.4 73.2±0.4 69.2±0.4 79.8±0.3 77.2±0.5 72.1±0.3 55.7±0.6 54.1±0.3 48.9±0.4
${\rm STABLE_{G}}$ 79.1±0.4 77.3±0.4 71.6±0.3 76.5±0.2 74.2±0.2 71.0±0.5 80.5±0.4 78.0±0.3 76.5±0.4 56.7±0.4 54.9±0.5 50.5±0.4
Ours 80.3±0.4 79.1±0.7 74.0±0.4 77.2±0.3 76.2±0.3 74.1±0.4 81.7±0.3 80.8±0.3 79.0±0.3 57.9±0.3 55.5±0.3 51.8±0.3
Dataset ($r$) Flickr (0.5%) Reddit (0.1%) Squirrel (1.7%) Gamers (0.5%)
Noise level (%) 5 10 25 5 10 25 5 10 25 5 10 25
Whole Dataset 48.8±0.1 47.0±0.1 46.4±0.1 93.4±0.0 92.7±0.1 91.9±0.0 39.8±0.6 34.5±0.7 31.5±0.4 59.0±0.1 58.4±0.1 58.0±0.1
GCond Plain 46.9±0.2 44.5±0.0 43.1±0.1 88.9±0.4 88.1±0.5 86.8±0.2 32.3±1.8 25.8±1.8 25.2±1.6 54.9±0.1 54.2±0.2 53.9±0.1
Jaccard 46.9±0.1 44.6±0.3 43.2±0.3 89.1±0.3 88.3±0.4 87.0±0.2 32.4±1.5 26.0±1.2 25.8±1.5 55.2±0.2 54.6±0.1 54.2±0.2
SVD 47.0±0.0 44.6±0.1 43.3±0.0 89.3±0.4 88.4±0.4 87.2±0.3 32.5±1.1 25.9±1.3 25.6±1.2 55.4±0.1 54.7±0.1 54.3±0.0
kNN 47.1±0.1 44.7±0.2 43.2±0.1 89.4±0.2 88.7±0.2 87.4±0.3 32.5±1.6 26.1±1.1 25.9±1.5 55.5±0.1 54.9±0.1 54.5±0.2
${\rm STABLE_{G}}$ 47.1±0.2 44.7±0.2 43.2±0.0 89.6±0.4 88.8±0.3 87.5±0.2 32.4±1.2 26.4±1.3 26.2±1.4 55.4±0.1 54.8±0.2 54.4±0.1
Ours 47.8±0.0 45.5±0.2 44.1±0.1 90.5±0.6 89.6±0.3 88.7±0.1 32.9±1.3 26.9±1.1 26.9±1.5 56.0±0.2 55.7±0.1 55.4±0.0
GCDM Plain 46.0±0.2 44.3±0.2 42.7±0.4 89.2±0.4 88.4±0.5 87.0±0.2 32.0±0.9 25.5±1.2 25.0±1.2 54.6±0.2 54.0±0.1 53.8±0.1
Jaccard 46.3±0.1 44.5±0.0 42.9±0.2 89.4±0.3 88.6±0.2 87.3±0.2 32.2±1.2 25.7±1.0 25.5±1.3 55.0±0.2 54.3±0.1 54.2±0.1
SVD 46.1±0.1 44.4±0.1 43.1±0.0 89.6±0.1 88.7±0.3 87.6±0.3 32.4±1.0 25.6±1.1 25.2±1.7 55.1±0.1 54.4±0.1 54.2±0.1
kNN 46.3±0.0 44.5±0.2 43.2±0.2 89.6±0.3 88.7±0.2 87.7±0.3 32.2±1.1 25.7±1.2 25.3±1.2 55.1±0.3 54.6±0.1 54.4±0.3
${\rm STABLE_{G}}$ 46.4±0.1 44.6±0.0 43.2±0.1 89.8±0.3 88.9±0.1 87.7±0.2 32.5±1.6 26.0±1.0 25.6±1.5 55.3±0.1 54.6±0.2 54.5±0.1
Ours 47.2±0.1 45.3±0.1 43.9±0.2 91.3±0.5 89.9±0.3 88.8±0.1 33.9±1.0 26.6±1.1 26.4±1.6 55.9±0.0 55.4±0.1 55.4±0.0
GDEM Plain 47.3±0.1 45.0±0.1 43.6±0.2 90.6±0.3 89.8±0.5 89.1±0.2 33.0±0.9 27.9±0.8 25.8±1.1 56.1±0.1 55.9±0.2 55.6±0.1
Jaccard 47.4±0.1 45.2±0.1 43.7±0.1 90.9±0.2 90.0±0.3 89.3±0.2 33.1±1.3 28.3±1.1 26.5±1.4 56.2±0.2 56.0±0.1 55.9±0.2
SVD 47.6±0.2 45.3±0.0 43.6±0.1 90.7±0.1 90.1±0.3 89.4±0.3 33.3±1.2 28.2±1.3 26.3±1.3 56.4±0.1 56.1±0.2 56.0±0.1
kNN 47.4±0.4 45.2±0.1 43.7±0.5 91.1±0.4 90.3±0.2 89.7±0.4 33.4±1.4 28.4±1.2 26.6±1.2 56.5±0.1 56.2±0.1 56.1±0.1
${\rm STABLE_{G}}$ 47.6±0.2 45.3±0.2 43.8±0.0 91.0±0.2 90.2±0.3 89.6±0.1 33.4±1.7 28.5±1.2 26.7±1.1 56.4±0.1 56.2±0.2 56.2±0.1
Ours 48.3±0.1 46.1±0.0 44.5±0.1 91.9±0.3 91.3±0.2 90.1±0.2 33.9±1.3 29.1±1.2 27.4±1.0 57.2±0.2 57.0±0.1 56.9±0.0
@@
Baselines and Evaluation Settings. We deploy our proposed method on three GC methods with different optimization strategies: (1) GCond [11]: the first GC method that matches the model gradients derived from the training and condensed graph; (2) GCDM [27]: an efficient GC method that minimizes the discrepancy between distributions of the training and condensed graphs. (3) GDEM [29]: a generalized GC method that aligns the training and condensed graph’s eigenvectors.
@@
Our primary objective is to enhance graph quality rather than merely optimizing model performance. Consequently, we select four state-of-the-art (SOTA) denoising techniques capable of addressing noise in inductive test settings. These methods are deployed with the aforementioned GC methods for comparison: (1) Jaccard [40]: it computes the Jaccard similarity for every pair of connected nodes and discards edges whose similarity is smaller than a specific threshold. (2) SVD [40]: it decomposes the noisy graph through Singular Value Decomposition (SVD) and adopts the low-rank approximation for the denoised graph to reduce the influence of potential high-rank noises. (3) kNN [41]: it identifies the k-nearest neighbors for each node based on their features and incorporates potentially effective edges into the original graph. (4) ${\rm STABLE_{G}}$ [25]: it refines the graph structure according to the node embeddings obtained from self-supervised training.
@@
Following [40], we assess our method’s resilience to both random noise and adversarial attacks. For random noise, it involves randomly adding fake edges to and removing real edges from the graph. Given the computational intensity of most adversarial attack methods, we utilize the SOTA scalable attack method, PR-BCD [22], to execute global adversarial attacks. We evaluate each method across different noise levels, defined by the ratio of changed edges in the original graph [40]. We evaluate random noise at levels {0%, 20%, 40%, 100%} and adversarial attacks at {5%, 10%, 25%}. We operate under the premise of uniform noise levels in real-world systems, indicating both training and test graphs share the identical noise level.